Grade 11Electric Current

Two trains A and B of length 400m each are moving on two parallel tracks with a uniform speed of 72km/h in the same direction, with A ahead of B. the driver of B decides to overtake A and accelerates by 1m/s2. After 50s, the end point of B moves past the head light of train A. what was the original distance between the two trains.

Radhika Batra
12 Years agoGrade 11

1 Answer

Premanshu Rao
11 Years ago
If the guard at the rear of B is passing the driver at the front of A then B has not only caught up the initial distance between the two trains but has also completely overtaken train A.  
Let the original separation distance be x. Then distance to be covered starting with relative velocity zero is s = 800 + x (800m being the overtaking distance from front meets rear to rear leaves front). 
Distance s covered at acceleration a starting from (relative) speed zero = 1/2 at^2 which for t=50 gives: 
= 0.5 x 50^2 = 1250m 
So 800 + x = 1250 Therefore x = 1250 -800 = 450m.